forgetmylife0:

A man is observing a pole of height 55 foot. According to his measurement, pole cast a 23 feet long shadow. Can you help him to know the angle of elevation of the sun from the tip of shadow?

Dva:

Dva:

tan x = 55/23 which is?

forgetmylife0:

2.391?

snowflake0531:

No, it's \(tan(x) = \frac{55}{23}\) so the answer would be \(tan^{-1}(\frac{55}{23})\)